OpenStudy (anonymous):

Write a polynomial function that has a zero at 2 and 1+i Attempt at solution

OpenStudy (anonymous):

(x+1+i)(x-1-i) x^2 - 2i (x^2-2i)(x-2) <--This will lead my no where!

OpenStudy (anonymous):

The answer is x^3 -4x^2+6x-4 if that could possible help

OpenStudy (anonymous):

forever alone -_-

OpenStudy (anonymous):

U guys are racist!

OpenStudy (anonymous):

*cries*

OpenStudy (anonymous):

*sobs*

OpenStudy (anonymous):

*stabs self with knife*

OpenStudy (anonymous):

will prevent bleeding if helped otherwise death -_-

OpenStudy (anonymous):

*bled to death*

OpenStudy (anonymous):

*comes back from dead*

OpenStudy (anonymous):

*I WILL HAUNT U ALL IF YOU DO NOT ANSWER MY QUESTION BOHOOOHOHOHOHOHOHOH FEAR ME!*

OpenStudy (anonymous):

*STABS GHOSTLY SELF IN CHEST ONCE MORE DUE TO A LACK OF ANSWERS*

OpenStudy (anonymous):

Hello my name is Steve Jobs I have stolen this guys soul and I have come back from the dead.

OpenStudy (anonymous):

To give this brown guy some pity for his suffering, i would appreciate it if someone answered his question.

OpenStudy (anonymous):

OK IF NO ONE ANSWERS HIS SIMPLE QUESTION I WILL PRESS THE SELF DESTRUCT SWITCH ON ALL THE APPLE DEVICES THAT I HAVE MADE

OpenStudy (anonymous):

&^#@&^$#&42*^(@#&(@*&)#(*)@# - THIS IS PROFANITY

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

what's this all about?

OpenStudy (anonymous):

^^^ Write a polynomial function that has a zero at 2 and 1+i

OpenStudy (anonymous):

It seems that you realized that if 1+i is a root, then its complex conjugate must also be a root. The complex conjugate of 1+i is 1-i, so you should start first by multiplying:\[(x-(1+i))(x-(1-i))=(x-1-i)(x-1+i)\]

OpenStudy (anonymous):

Everything else you did was correct, you just picked the wrong things to multiply in the very first step.

OpenStudy (anonymous):

Ohhh thank you, I really appreciate it :)

OpenStudy (anonymous):

*Gives joemath314159 <--- Hey THATS PI ! YUM 5 KARMA* *Takes 14 KARMA away from everyone else*

OpenStudy (anonymous):

If \(r_1,\dots r_n\) are the roots of a polynomial \(P\), we have that: \[P(x)=(x-r_1)\cdots(x-r_n)\] If \(r_1 \in \mathbb{C}\), then \(\bar{r_1}\in \mathbb{C}\) is also a root. So, while the problem just says 2 and 1+i, 1-i is also included as a root. So, you have: \[\begin{align} P(x)&=(x-2)(x-(1+i))(x-(1-i))\\&=(x-2)(x-1-i)(x+1+i)\\ &=. . . \end{align}\]

OpenStudy (anonymous):

Thanks for helping aswell Ive got my answer and I think getting there will stick due to this episode. I cant believe people would help out a brown guys. *Hugs limitless & Joemath*